# Machine Learning $H$-theorem

### v1

@@ I Introduction
The discovery of $H$-theorem by Ludwig Boltzmann in 1872 boltzmann1872waermegleichgewicht played a pivotal role in establishing statistical mechanics as a framework that provides thermodynamics with a microscopic origin and liberates it from its axiomatic foundations. At the same time, $H$-theorem was for a long time the topic of much debate https://doi.org/10.1002/andp.18972960216 ; zermelo2 ; zermelo1 , parts of which persist to this day Villani2003 ; Borsoni2024 ; https://doi.org/10.1002/cpa.10012 ; DiPerna1989 . One lingering issue is the Loschmidt paradox Cercignani2006-if ; Cercignani1988 , which relates to the remarkable feature of $H$-theorem that there exists an $H$-functional which can only decrease in time and therefore signal the presence of an arrow of time eddington1929nature despite the system being microscopically reversible doi:10.1073/pnas.11.7.436 , which means that the particles of which the gas is composed undergo interactions that are time-reversal invariant. This paradox is now better understood, in part because of the discovery of BBGKY hierarchy grad1949 ; ecc82589-c02d-30e9-9891-fe9629a2fd13 , which must be truncated by assuming molecular chaos so that one is left with the Boltzmann equation. The assumption of molecular chaos introduces an arrow of time; while Lanford’s theorem lanford1975 proves its validity in the Boltzmann–Grad limit grad1949 for short times, no general proof exists for arbitrary times or interaction types, so its adoption remains, to some extent, a leap of faith. Nevertheless, the Boltzmann equation has been found to be a valuable tool for deriving continuum equations from first principles that has been very widely used for more than a century Cercignani1988 ; chapman1990mathematical in fields ranging from solid state physics PhysRevB.97.045105 ; mueller2009graphenenearlyperfect ; PhysRevLett.115.216806 to high energy physics De_Groot1980-px ; Cercignani2002-lo ; Mueller_2004 as well as in active matter PhysRevE.83.030901 ; PhysRevE.74.022101 ; grosvenor2025hydrodynamicsboostinvariancekinetictheory ; rubenliermicroscopic .
@@
In this work, we learn to output a value that resembles $H$-functional up to an affine transformation for randomly initialized hard disks PhysRevA.1.18 ; PhysRevE.92.022131 ; PhysRevLett.18.988 ; ORBAN1967620 . We do this using an approach that is similar to that of Ref. rahaman2019learningarrowtime where a linear function is derived whose cost function is such that the learning processes pushes the model to learn to attribute higher output values to input values corresponding to later times.
@@ II Boltzmann equation
We consider the number of particles $dN$ at a region in phase space to be given by
@@
$\begin{split}&\left(\frac{\partial f}{\partial t}\right)_{\text{self}}=\\ &\iint d\mathbf{v}_{2}db\,u\left(f(\mathbf{v},t)f(\mathbf{v}_{2},t)-f(\mathbf{v}^{\prime}_{1},t)f(\mathbf{v}^{\prime}_{2},t)\right),\end{split}$ (3)
@@ III Hard disk simulation
$\mathbf{n}\cdot\mathbf{u}=-\mathbf{n}\cdot\mathbf{u}^{\prime}~~,~~\mathbf{z}\times\mathbf{n}\times\mathbf{u}=\mathbf{z}\times\mathbf{n}\times\mathbf{u}^{\prime}~~,$ (7)
@@
$\mathbf{u}^{\prime}\cdot\mathbf{n}<0~~.$ (8)
as otherwise the particles are moving away from each other. If this constraint is not imposed, particles will get stuck together. The radius of the disks is taken to be $d=3\cdot 10^{-3}$ whereas the periodic box has a unit surface. When the frame rate is too small, it is possible that certain collisions are missed. In Fig. 2, we show that at a frame rate of around $60$, the number of collisions starts to plateau, so this is the frame rate we take.
@@ IV Machine learning $H$-theorem
Let us now describe the machine learning process. At every training step, we have as input sequential data $V^{[n]}_{t}$ for all $t$ and $n$, where $t$ corresponds to the time step and $n$ corresponds to the run. For any run $n$, we have
$\mathbf{v}^{(i),[n]}_{t}\in V^{[n]}_{t}~~$, (9)
@@
$\begin{split}&\mathcal{L}^{(0)}\left(V^{[n]}_{t},V^{[n]}_{t+1}\right)=h\left(V^{[n]}_{t+1}\right)-h\left(V^{[n]}_{t}\right)\\ &+\lambda\,\!\left(h\left(V^{[n]}_{t+1}\right)-h\left(V^{[n]}_{t}\right)\right)^{2}\end{split}$ (10b)
@@
After experimenting with (10), it is found that the trained function $h(V_{t})$, unlike the real $H$-functional, displays strongly oscillatory behavior upon increasing the time step. To mitigate this, we note that (6a) is an inequality which only requires that the $H$-functional does not increase as a function of time, not that it decreases. Therefore, we consider a generalization of the loss function of (10b) which is given by
$\begin{split}&\mathcal{L}^{\alpha}\left(V^{[n]}_{t},V^{[n]}_{t+1}\right)\\ =&-\text{leaky-ReLu}\left(h\left(V^{[n]}_{t}\right)-h\left(V^{[n]}_{t+1}\right),\alpha\right)\\ &+\lambda\,\!\left(h\left(V^{[n]}_{t+1}\right)-h\left(V^{[n]}_{t}\right)\right)^{2}~~.\end{split}$ (11)
It follows that when $0\leq\alpha<1$, (11) makes sure that the function $h(V_{t})$ is punished more for incorrectly outputting a lower value at a later time step than it is rewarded for correctly outputting a higher value at later times. This choice is not only consistent with (6a), which only imposes the former on the $H$-functional, but also turns out to reduce oscillation of $h(V_{t+1})$. We optimize for a total number of epochs $N_{e}=120$ with the loss of (11) averaged over runs $n$ and time steps $t$ using the Adam optimizer kingma2017adammethodstochasticoptimization , choosing learning rate $4\cdot 10^{-4}$ and $\lambda=0.2$. All the hidden layers of the model have hidden dimension $n_{\text{hidden}}$ and ReLu-activations (see Fig. 4 and Table 1). While training, we increase $\alpha$ as a function of epoch $e$ as
@@
Because of the choice of a gradual increase according to (12) with a high $\alpha_{\text{final}}$, one eventually reaches a point for all of the models where $\mathcal{L}_{t}^{\alpha(e)}>0$, which indicates that the trivial solution where $h(V_{t+1})-h(V_{t})=0$ for all $t$ is no longer suboptimal. This means the the model can no longer improve and the model output difference between time steps starts to gradually move to zero. We call the epoch for which this first happens the critical epoch $e_{\text{c}}$. Table (2) gives the average loss for the different models at the final epoch together with $e_{\text{c}}$.
@@
V Affine transformation
Having fully specified the learning process of the model $h(V_{t})$, let us consider its relation to the $H$-functional provided in (5). The model $h(V_{t})$ can only possibly resemble the $H$-functional up to an affine transformation kuroyanagi2025deeplearningthermodynamiclaws ; rahaman2019learningarrowtime . This is because (11) is invariant under shifts constant in time. Furthermore, we have for $c\in\mathbb{R}$
$\{h(V_{t}),\lambda\}\rightarrow\{ch(V_{t}),\lambda/c\}\Rightarrow\mathcal{L}^{\alpha}_{t}\rightarrow c\mathcal{L}^{\alpha}_{t}~~,$ (14)
which means we can multiply the function $h(V_{t})$ with a constant factor and it will only affect the loss with a total factor provided we rescale the physically meaningless $\lambda$. Therefore, in order to compare the model to the real $H$-functional we transform the function $h(V_{t})$ as
$\tilde{h}^{[n]}\left(V^{[n]}_{t}\right)=a^{[n]}h\left(V^{[n]}_{t}\right)+b^{[n]}~~.$ (15)
@@ V Affine transformation
$\operatorname*{argmin}_{a^{[n]},b^{[n]}}=\sum_{t}\left(\tilde{h}^{[n]}\left(V^{[n]}_{t}\right)-H_{t}^{[n]}\right)^{2}~~.$ (16)
In Fig. 5, the result is shown for the different models and compared to the $H$-functional. It is clear that for $n_{\text{hidden}}=32$ the model performs better than for $n_{\text{hidden}}=16$, whereas there is no big difference between $n_{\text{hidden}}=64$ ad $n_{\text{hidden}}=32$. Although the affine parameters are fitted for each run separately, we would ideally have a model that can provide a value for the $H$-functional that is consistent throughout all runs. To find out whether the model is consistent throughout runs, we compute the average and the variance of the affine parameters in Table 3. We see that going from $n_{\text{hidden}}=16$ to $n_{\text{hidden}}=32$ lowers the error, which indicates a better learning process, whereas $n_{\text{hidden}}=64$ slightly increases it, although the coefficient of variation of the affine parameters is lowered significantly, suggesting that it is a more coherent model. Note that $\langle a\rangle$ is always negative as the model learns a function that increases in time whereas the $H$-functional decreases in time as required by (6a).
@@ VI Discussion
In this work we used a siamese neural network to learn the $H$-functional of $H$-theorem for sequential data of a thousand hard disks colliding with each after being initialized in a state of low entropy. The loss function that was formulated in order to learn $H$-theorem was based on the requirement that the learned function should attribute higher values to input data corresponding to later points in time. This requirement forces the model to identify the arrow of time which according to $H$-theorem was predicted to exist for this simple but also infinitely chaotic model of hard disks. The resulting model only resembles the extracted $H$-functional after an appropriately fitted affine transformation. Two things that were found to be important for making the model resemble the $H$-functional are
@@
For future work, it would be interesting to explore more complicated types of collisional interactions, such as for example polyatomic gases Cercignani1981 or for active gas of two species of hard spheres that collide reactively and can resemble the phenomenon of bird flocking rubenliermicroscopic .

### v3

@@ I Introduction
The discovery of $H$-theorem by Ludwig Boltzmann in 1872 boltzmann1872waermegleichgewicht played a pivotal role in establishing statistical mechanics as a framework that provides thermodynamics with a microscopic origin and liberates it from its axiomatic foundations. At the same time, $H$-theorem was for a long time the topic of much debate https://doi.org/10.1002/andp.18972960216 ; zermelo2 ; zermelo1 , parts of which persist to this day Villani2003 ; Borsoni2024 ; https://doi.org/10.1002/cpa.10012 ; DiPerna1989 . One lingering issue is the Loschmidt paradox Cercignani2006-if ; Cercignani1988 , which relates to the remarkable feature of $H$-theorem that there exists an $H$-functional which can only decrease in time and therefore signal the presence of an arrow of time eddington1929nature despite the system being microscopically reversible doi:10.1073/pnas.11.7.436 , which means that the particles of which the gas is composed undergo interactions that are time-reversal invariant. This paradox is now better understood, in part because of the discovery of BBGKY hierarchy grad1949 ; ecc82589-c02d-30e9-9891-fe9629a2fd13 , which must be truncated by assuming molecular chaos so that one is left with the Boltzmann equation. The assumption of molecular chaos introduces an arrow of time; while Lanford’s theorem lanford1975 ; AST_1976__40__117_0 ; RevModPhys.52.569 proves its validity in the Boltzmann–Grad limit grad1949 for short times, no general proof exists for arbitrary times or interaction types, so its adoption remains, to some extent, a leap of faith. Nevertheless, the Boltzmann equation has been found to be a valuable tool for deriving continuum equations from first principles that has been very widely used for more than a century Cercignani1988 ; chapman1990mathematical in fields ranging from solid state physics PhysRevB.97.045105 ; mueller2009graphenenearlyperfect ; PhysRevLett.115.216806 to high energy physics De_Groot1980-px ; Cercignani2002-lo ; Mueller_2004 as well as in active matter PhysRevE.83.030901 ; PhysRevE.74.022101 ; grosvenor2025hydrodynamicsboostinvariancekinetictheory ; rubenliermicroscopic .
@@
In this work, we learn to output a value that resembles $H$-functional up to an affine transformation for sequential data of a numerical simulation of randomly initialized hard disks PhysRevA.1.18 ; PhysRevE.92.022131 ; PhysRevLett.18.988 ; ORBAN1967620 . We infer the $H$-functional using an approach that is similar to that of Ref. rahaman2019learningarrowtime where a linear function is derived whose cost function is such that the learning processes pushes the model to learn to attribute higher output values to input values corresponding to later times.
@@ II Boltzmann equation
We consider the number of particles $dN$ in a region in phase space to be given by
@@
$\begin{split}&\left(\frac{\partial f}{\partial t}\right)_{\text{self}}=\\ &\iint d\mathbf{v}_{2}db\,u\left(f(\mathbf{v}^{\prime}_{1},t)f(\mathbf{v}^{\prime}_{2},t)-f(\mathbf{v},t)f(\mathbf{v}_{2},t)\right),\end{split}$ (3)
@@ III Hard disk simulation
$\mathbf{n}\cdot\mathbf{u}^{\prime}=-\mathbf{n}\cdot\mathbf{u}~~,~~\mathbf{z}\times\mathbf{n}\times\mathbf{u}^{\prime}=\mathbf{z}\times\mathbf{n}\times\mathbf{u}~~,$ (7)
@@
$\mathbf{u}\cdot\mathbf{n}<0~~.$ (8)
as otherwise the particles are moving away from each other. If this constraint is not imposed, particles will get stuck together. The diameter of the disks is taken to be $d=6\cdot 10^{-3}$ whereas the periodic box has a unit surface. When the frame rate is too small, it is possible that certain collisions are missed. In Fig. 2, we show that at a frame rate of around $60$, the number of collisions starts to plateau, so this is the frame rate we take.
@@ IV Machine learning $H$-theorem
Let us now describe the machine learning process. At every training step, we have as input sequential data $\mathbf{V}^{[n]}_{t}$ for all $t$ and $n$, where $t$ corresponds to the time step and $n$ corresponds to the run. For any run $n$, we have
$\mathbf{v}^{(i),[n]}_{t}\in\mathbf{V}^{[n]}_{t}~~$, (9)
@@
$\begin{split}&\mathcal{L}^{(0)}\left(\mathbf{V}^{[n]}_{t},\mathbf{V}^{[n]}_{t+1}\right)=h\left(\mathbf{V}^{[n]}_{t}\right)-h\left(\mathbf{V}^{[n]}_{t+1}\right)\\ &+\lambda\,\!\left(h\left(\mathbf{V}^{[n]}_{t+1}\right)-h\left(\mathbf{V}^{[n]}_{t}\right)\right)^{2}\end{split}$ (10b)
@@
After experimenting with (10), it is found that the trained function $h(\mathbf{V}_{t})$, unlike the real $H$-functional, displays strongly oscillatory behavior upon increasing the time step. To mitigate this, we note that (6a) is an inequality which only requires that the $H$-functional does not increase as a function of time, not that it decreases. Therefore, we consider a generalization of the loss function of (10b) which is given by
$\begin{split}&\mathcal{L}^{\alpha}\left(\mathbf{V}^{[n]}_{t},\mathbf{V}^{[n]}_{t+1}\right)\\ =&\text{leaky-ReLu}\left(h\left(\mathbf{V}^{[n]}_{t}\right)-h\left(\mathbf{V}^{[n]}_{t+1}\right),\alpha\right)\\ &+\lambda\,\!\left(h\left(\mathbf{V}^{[n]}_{t+1}\right)-h\left(\mathbf{V}^{[n]}_{t}\right)\right)^{2}~~.\end{split}$ (11)
It follows that when $0\leq\alpha<1$, (11) makes sure that the function $h(\mathbf{V}_{t})$ is punished more for incorrectly outputting a lower value at a later time step than it is rewarded for correctly outputting a higher value at later times. This choice is not only consistent with (6a), which only imposes the former on the $H$-functional, but also turns out to reduce oscillation of $h(\mathbf{V}_{t+1})$. We optimize for a total number of epochs $N_{e}=120$ with the loss of (11) averaged over runs $n$ and time steps $t$ using the Adam optimizer kingma2017adammethodstochasticoptimization , choosing learning rate $4\cdot 10^{-4}$ and $\lambda=0.2$. All the hidden layers of the model have hidden dimension $n_{\text{hidden}}$ and ReLu-activations (see Fig. 4 and Table 1). While training, we increase $\alpha$ as a function of epoch $e$ as
@@
Because of the choice of a gradual increase according to (12) with a high $\alpha_{\text{final}}$, one eventually reaches a point for all of the models where $\mathcal{L}_{t}^{\alpha(e)}>0$, which indicates that the trivial solution where $h(\mathbf{V}_{t+1})-h(\mathbf{V}_{t})=0$ for all $t$ is no longer suboptimal. This means the model can no longer improve and the model output difference between time steps starts to gradually move to zero. We call the epoch for which this first happens the critical epoch $e_{\text{c}}$. Table (2) gives the average loss for the different models at the final epoch together with $e_{\text{c}}$.
@@
V Comparison with $H$-functional
Having fully specified the learning process of the model $h(\mathbf{V}_{t})$, let us consider its relation to the $H$-functional provided in (5). The model $h(\mathbf{V}_{t})$ can only possibly resemble the $H$-functional up to an affine transformation kuroyanagi2025deeplearningthermodynamiclaws ; rahaman2019learningarrowtime . This is because (11) is invariant under shifts constant in time. Furthermore, we have for $c\in\mathbb{R}$
$\{h(\mathbf{V}_{t}),\lambda\}\rightarrow\{ch(\mathbf{V}_{t}),\lambda/c\}\Rightarrow\mathcal{L}^{\alpha}_{t}\rightarrow c\mathcal{L}^{\alpha}_{t}~~,$ (14)
which means we can multiply the function $h(\mathbf{V}_{t})$ with a constant factor and it will only affect the loss with a total factor provided we rescale the physically meaningless $\lambda$. Therefore, in order to compare the model to the real $H$-functional we transform the function $h(\mathbf{V}_{t})$ as
$\tilde{h}^{[n]}\left(\mathbf{V}^{[n]}_{t}\right)=a^{[n]}h\left(\mathbf{V}^{[n]}_{t}\right)+b^{[n]}~~.$ (15)
@@ V Comparison with $H$-functional
$\operatorname*{argmin}_{a^{[n]},b^{[n]}}=\sum_{t}\left(\tilde{h}^{[n]}\left(\mathbf{V}^{[n]}_{t}\right)-H_{t}^{[n]}\right)^{2}~~.$ (16)
In Fig. 5, the result is shown for the different models and compared to the $H$-functional. It is clear that for $n_{\text{hidden}}=32$ the model performs better than for $n_{\text{hidden}}=16$, whereas there is no big difference between $n_{\text{hidden}}=64$ ad $n_{\text{hidden}}=32$. Although the affine parameters are fitted for each run separately, we would ideally have a model that can provide a value for the $H$-functional that is consistent throughout all runs. To find out whether the model is consistent throughout runs, we compute the average and the variance of the affine parameters in Table 3. We see that going from $n_{\text{hidden}}=16$ to $n_{\text{hidden}}=32$ lowers the error, which indicates a better learning process, whereas $n_{\text{hidden}}=64$ slightly increases it, although the coefficient of variation of the affine parameters is lowered significantly, suggesting that it is a more coherent model. Note that $\langle a\rangle$ is always negative as the model learns a function that increases in time whereas the $H$-functional is defined to be decreasing in time as follows from (6a).
@@ VI Discussion
In this work we used a siamese neural network to learn the $H$-functional of $H$-theorem for sequential data of a numerical simulation of a thousand hard disks colliding with each other after being initialized in a low-entropy state. The loss function that was formulated in order to learn $H$-theorem was based on the requirement that the learned function should attribute higher values to input data corresponding to later points in time. This requirement forces the model to identify the arrow of time which according to $H$-theorem was predicted to exist for this simple but also infinitely chaotic model of hard disks. The resulting model only resembles the extracted $H$-functional after an appropriately fitted affine transformation. Two things that were found to be important for making the model resemble the $H$-functional are
@@
In future work, it would be interesting to explore more complicated types of collisional interactions, such as for example polyatomic gases Cercignani1981 or for active gas of two species of hard spheres that collide reactively and can resemble the phenomenon of bird flocking rubenliermicroscopic .